# Enumerate–Conjecture–Prove: Formally Solving Answer-Construction Problems in Math Competitions

### v2

@@ 4.1 Data Sources and Scope
The dataset spans a wide difficulty spectrum, ranging from challenging high-school-level problems to sophisticated Olympiad-style questions. It covers diverse mathematical domains including Algebra, Number Theory, Combinatorics, Probability, and Calculus. Rather than explicit domain labeling, we employ structural filtering to exclusively retain answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
@@ 4.2 Data Deduplication
The dataset initially aggregated 5,555 problems. To remove duplication, both string-based matching and embedding-based duplicate filtering methods were utilized, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold. This process resulted in 5,316 unique and diverse problems.
@@ 4.3 Problem Autoformalization
We utilized our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into formally stated Lean theorems. Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with gpt-4o model serving as an LLM judge to verify semantic consistency.
Each problem went up to $T=5$ refinement iterations per model. This procedure yielded a well-formed subset of 3,431 problems, whose formalizations are compilable in the Lean compiler, passed LLM-based semantic check, and formatted according to the template. Each entry of ConstructiveBench includes the informal and formal problem statement and ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty. Figure 4 illustrates an example dataset entry alongside the distribution of answer types present in ConstructiveBench. The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various mathematical types of problems.
Furthermore, to address potential data contamination, where LLMs might inadvertently have memorized training examples, we explicitly include a carefully selected test subset of 92 problems from the formalized data created after June 2024, beyond the knowledge cutoff date of most LLMs, including the ones that will be evaluated (gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
@@ 5 Empirical Evaluation
    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success on answer-construction problems compared to purely symbolic or LLM-only approaches?
@@ 5.1 Experimental Setup
Answer-Construction
For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically-correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers). We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, and deepseek-v3. The models are given the formalized lean statement, and the type of answer to be constructed. We compare the ECP framework against baseline Chain-of-Thought (CoT) method (i.e. without enumerating and pattern finding) for answer construction. The constructed answer is considered correct iff the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics simp, aesop, nlinarith, ring, norm_num).
Theorem-Proving
We evaluate these theorem-proving tasks using state-of-the-art theorem provers, comparing baseline symbolic methods and specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B. Evaluations are performed with one A100 GPU and 64 CPU cores for Lean verification, using hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048 and Pass@@@@32 metric. The experiment includes 2 conjecturing setups (CoT or ECP), 3 different models for conjecturing, and 3 different prover models for proving, resulting in a total of 18 experiments. We use 64 AMD EPYC 7453 processor CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours using an A100 80GB GPU for 2 hours per setup.
@@ 5.2 Results
Autoformalization
Out of the 5,316 problems in the ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks and, therefore, yielding an overall success rate of 64.6%. The information retriever contributed 18,516 relevant Lean definitions and theorems based on error messages, helping recover 569 additional formalizations.
Answer-Construction Accuracy
We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified problems from the ConstructiveBench and the subset of 337 problems that require answer construction from 640 problems in PutnamBench. Answer correctness was rigorously verified through formal proofs in Lean.
@@
Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and 337 PutnamBench answer-construction problems between CoT baseline and ECP-assisted methods.
5.3 Theorem-Proving Accuracies
We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem-proving using the same ConstructiveBench problems. Constructed answers generated in Section 5.2 are substituted back into the formal statements, reducing answer-construction tasks into standard theorem-proving tasks.
@@ 5.3 Theorem-Proving Accuracies
Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimini-Prover-7B-Preview DeepSeek-Prover-V2-7B
@@
Table 3: Theorem proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
The summarized results on the ConstructiveBench problems are presented in Table 3. Among evaluated provers, DeepSeek-Prover-V2-7B consistently exhibited the best proving capability within the ECP framework, achieving the highest proving accuracy of 25.01% with gpt-4.1-mini. Goedel-Prover-SFT and Kimina-Prover-7B-Preview showed substantial improvements from baseline to ECP as well, though slightly behind DeepSeek-Prover-V2-7B. Additionally, on the subset of 337 problems requiring answer construction from PutnamBench, by using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, outperforming to 0.59% accuracy with CoT for answer construction. Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model-prover combinations.
@@ 6 Limitations and Future Work
Quality of Autoformalized Theorem Statements. Our framework and dataset depend heavily on the quality of autoformalization produced by black-box LLMs, which are often trained on limited aligned informal-formal mathematical data (Wu et al., 2022b). In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics. While manual evaluation offers high fidelity, it is time-consuming and difficult to scale. Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence to the original problem. Future work should explore improved evaluation metrics for formalization quality, as well as expanding datasets that better align informal and formal representations of mathematical problems.
Theorem-proving: Our proving stage uses the off-the-shelf theorem provers, which leaves significant room for optimization. In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could lead to improvements in success rates and runtime. Another limitation is the combinatorial explosion during the enumeration phases for difficult IMO problems, where the time complexity of the enumeration program blows up as the in-context parameter increases. Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space. Fine-tuning LLMs or developing search-based algorithms on the formal answer construction task thus represents a promising avenue for next-generation mathematical AI models.
@@ 7 Conclusion
We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean. To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedbacks. In empirical evaluations, ECP improved answer-construction accuracy with gpt-4.1-mini from 14.54% under direct CoT prompting to 45.06%. Moreover, when paired with DeepSeek-Prover-V2-7B, ECP formally proved 858 of 3,431 problems (25.01%) end-to-end. These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer construction mathematics problems. Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to challenging Olympiad and research-level mathematical problems.

### v3

@@ 4.1 Data Sources and Scope
The dataset spans a wide difficulty spectrum, ranging from challenging high-school-level problems to sophisticated Olympiad-style questions. It covers diverse mathematical domains including algebra, number theory, combinatorics, probability, and calculus. Rather than explicit domain labeling, we employ structural filtering to retain exclusively answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
@@ 4.2 Data Deduplication
The dataset initially aggregated 5,555 problems. To remove duplication, we utilized both string-based matching and embedding-based duplicate filtering methods, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold. This process resulted in 5,316 unique and diverse problems.
@@ 4.3 Problem Autoformalization
We used our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into a formally stated Lean theorem. Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with the gpt-4o model serving as an LLM judge to verify semantic consistency.
Each problem underwent up to $T=5$ refinement iterations per model. This procedure yielded a well-formed subset of 3,431 problems, whose formalizations compile successfully in Lean, pass the LLM-based semantic check, and are formatted according to the template. Each entry in ConstructiveBench includes the informal and formal problem statements, the ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty. Figure 4 illustrates an example dataset entry alongside the distribution of answer types present in ConstructiveBench. The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various types of mathematical problems.
Furthermore, to address potential data contamination—where LLMs might inadvertently have memorized training examples—we explicitly include a carefully selected test subset of 92 problems drawn from formalized data created after June 2024, which is beyond the knowledge cutoff date of most LLMs (including those we will evaluate: gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
@@ 5 Empirical Evaluation
    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success rates on answer-construction problems compared to purely symbolic or LLM-only approaches?
@@ 5.1 Experimental Setup
Answer-Construction.
For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers). We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, as well as deepseek-v3. The models are given the formalized Lean statement and the type of answer to be constructed. We compare the ECP framework against a baseline Chain-of-Thought (CoT) method (i.e., without enumeration and pattern finding) for answer construction. The constructed answer is considered correct if and only if the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics: simp, aesop, nlinarith, ring, norm_num).
Theorem-Proving.
We evaluate the theorem-proving tasks using state-of-the-art provers, comparing baseline symbolic methods with specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B. Evaluations are performed on one A100 GPU and 64 CPU cores for Lean verification, using the hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048, and the Pass@32 metric. The experiment includes two conjecturing setups (CoT or ECP), three different models for conjecturing, and three different prover models for proving, resulting in a total of 18 experiments. We use 64 AMD EPYC 7453 CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours on an A100 80 GB GPU (2 hours per setup).
@@ 5.2 Results
Autoformalization.
Out of the 5,316 problems in ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks, yielding an overall success rate of 64.6%. The information retriever contributed 18,516 relevant Lean definitions and theorems based on error messages, helping recover 569 additional formalizations.
Answer-Construction.
We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified ConstructiveBench problems and on the subset of 337 answer-construction problems from the 640 problems in PutnamBench. Answer correctness was rigorously verified through formal proofs in Lean.
@@
Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and on 337 PutnamBench answer-construction problems.
Theorem-Proving.
We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem proving using the same ConstructiveBench problems. Constructed answers generated in Section 5.2 are substituted back into the formal statements, reducing answer-construction tasks into standard theorem-proving tasks.
@@
Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimina-Prover-7B-Preview DeepSeek-Prover-V2-7B
@@
Table 3: Theorem-proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
The summarized results on the ConstructiveBench problems are presented in Table 3. Among evaluated provers, DeepSeek-Prover-V2-7B consistently exhibited the best proving capability within the ECP framework, achieving the highest proving accuracy of 25.01% with gpt-4.1-mini. Goedel-Prover-SFT and Kimina-Prover-7B-Preview showed substantial improvements from baseline to ECP as well, though slightly behind DeepSeek-Prover-V2-7B. Additionally, on the subset of 337 answer-construction problems from PutnamBench, using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, compared to 0.59% with CoT for answer construction. Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model–prover combinations.
@@ 6 Limitations and Future Work
Quality of Autoformalized Theorem Statements. Our framework and dataset depend on the quality of autoformalization produced by black-box LLMs. In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics. While manual evaluation offers high fidelity, it is time-consuming and difficult to scale. Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence with the original problem. Future work should explore improved evaluation metrics for formalization quality and expand datasets that better align informal and formal representations of mathematical problems.
Theorem-Proving. Our proving stage uses off-the-shelf theorem provers, which leaves significant room for optimization. In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could improve both success rates and runtime. Another limitation is the combinatorial explosion during the enumeration phase for difficult IMO problems, where the time complexity of the enumeration program grows rapidly as the in-context parameter increases. Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space. Fine-tuning LLMs or developing search-based algorithms for the formal answer-construction task thus represents a promising avenue for next-generation mathematical AI models.
@@ 7 Conclusion
We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean. To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedback. In empirical evaluations, ECP improved answer-construction accuracy with gpt-4.1-mini from 14.54% under direct CoT prompting to 45.06%. Moreover, when paired with DeepSeek-Prover-V2-7B, ECP formally proved 858 of 3,431 problems (25.01%) end-to-end. These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer-construction mathematics problems. Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to challenging Olympiad and research-level mathematical problems.